if the degree measures of 2 comlementary angles are in the ratio 2:13, what is the measure of the smaller angle?
Answer by mathslover(157)

Sum of the complementary angles is 90 degress.
If one of the angles is 2x , the other angle is 13x
(from the given condition that their ratios are 2:13)
=> 2x + 13x = 90
=> 15x = 90
=> x=6
Measure of the smaller angle is 2x = 2* 6 =12 degrees